# Invisible Labor in Open Source Software Ecosystems

### v1

@@ 3. Methods
Within OSS ecosystems, manual and automated tools have had mixed success at measuring diverse activities, code and noncode alike. As an alternative, we used a participatory feminist technique: working with developers to design a survey, and then directly asking a global sampling of developers about their OSS activities. The following sections describe how we designed and conducted the survey.
@@ 3.1. Survey Instrument Design
We codesigned a Qualtrics web survey (Tourangeau et al., 2013) with practicing developer advocates (among the coauthors). Given the importances of visibility and compensation for invisible labor, we dedicated a portion of the survey to each. We iteratively developed the questions and their responses among ourselves and with our Institutional Review Board (IRB), which approved the design before the start of the study.
Broadly, the survey focused participant attention on recent collaborative OSS projects and the breadth of tasks performed therein. This gave us access to participant perceptions of tasks beyond code contributions while still including coding activities. We asked a series of questions about the credit they receive as one form of compensation. Then, we asked several questions about how often other people know about the tasks they perform to measure labor visibility. See Appendix A for the full survey.
@@ 3.1.1. Guiding Participant Attention
A participant’s experience with a survey begins with recruitment materials. These materials, along with the consent form (Q1) and other questions throughout the survey, needed to carefully avoid the term “invisible labor” because topics related to supportive labor tend to be femininely coded — and even denigrated — in many global contexts (Berdahl et al., 2018; D’Ignazio and Klein, 2020). Using the term “invisible labor” in a survey could incentivize gendered signaling of masculinity and femininity or political signaling. Instead, we devised other terms to help participants recall the full breadth of tasks that people perform in their projects. We advertised the survey around “receiving credit” for their work as this is an important concept to many in OSS communities (Young et al., 2021) but which should be less likely to evoke either self-selection into the survey based on the survey focus or gendered social desirability bias (Krumpal, 2013).
@@
The survey began by prompting the participant to focus on their experiences with open source projects that they worked on with other people during the last two years (Q2). We limited the survey to this time period to reduce experience underretrieval due to excessive cognitive load (Lenzner et al., 2010; Tourangeau, 2018). Focusing their attention on collaborative open source projects then gave us access to OSS projects that pose the greatest challenge to sustainability (Eghbal, 2020).
@@
Later, we focused the participant’s attention on how “tasks vary in shape, size, and how visible they are to other people” (Q5). We asked them to “think through the different kinds of tasks you have done as part of these projects over the past 2 years.” We did this to provide rich retrieval cues about varied tasks, which tend to improve recall (Tourangeau, 2018). This broadened their perception out from “just code” to include as many tasks as possible within their work, thereby bringing their visible and invisible labor to the fore. For those who already considered noncoding activities to be contributions, this would not affect their judgment; however, for those who might consider code the only legitimate type of “contribution” (Eghbal, 2020; Casari et al., 2021), this design avoided the term contribution (via “task” instead) while improving information retrieval about diverse tasks (Tourangeau, 2018).
@@ 3.1.2. Compensation Questions
This followed with several questions about receiving credit for tasks, which we defined for them as “the project told someone that you did something for the project” (Q6-10). We asked about how many projects they received credit from (via 5-point Likert scale), how many tasks they received credit for (Likert scale), the medium through which they received credit (multiple select), and their satisfaction with the credit they received (Likert scale). These questions helped us gauge how much compensation (here, attribution) individuals receive, how they receive it, and to measure any relationship strength between attribution and perceptions of visibility.
@@ 3.1.3. Labor Visibility Questions
Unlike in crowd work where all work might take place on one platform (Toxtli et al., 2021), OSS activities take place across an ecosystem (Casari et al., 2021) that makes measurement via tracking difficult. We therefore measured labor visibility by asking participants how often different numbers of people knew about tasks they performed. This approximated how visible different tasks are to OSS communities. Behavioral frequency impressions (e.g. how often someone performs invisible tasks) are prone to under- and overestimation (Conrad et al., 1998). While we cannot directly verify their perceptions, our approach treats their perceptions as expertise (Bardzell, 2010). Combining randomization with human cognitive biases allowed us to estimate likely bounds on how much labor individuals perform at different visibility levels.
To that end, we began the labor visibility questions with a prompt (“Think about all the tasks you performed again,” Q11) to aid recall of their activities. Then, we presented participants with three questions:
@@
Information presented earlier in a survey tends to produce assimilation effects — that is, later judgments are moved in the direction of earlier judgments (Tourangeau, 2018), reminiscent of the anchoring and adjustment heuristic or priming effects (Kahneman et al., 1982; Molden, 2014). So if we only presented Q12, Q13, and Q14 in that order (called $\geq 2$, 1, 0 hereafter for clarity), participant responses would have been likely to bias results toward presumptions of high visibility (2 or more people, or “openness”). Therefore, we randomized whether participants saw these questions in order of what we call anchoring to high visibility ($\geq 2$, 1, 0) or anchoring to low visibility (0, 1, $\geq 2$).
@@
Following these, we asked participants how important it is to receive credit for tasks (Q16) along with several free response questions about what did and did not work well for them about how people give credit.
3.2. Study Procedures
@@
To test the survey instrument and measure invisible labor in OSS, we recruited 142 participants between January and June, 2022. We did this by posting IRB-approved recruitment materials on the social media platform formerly known as Twitter with the hashtag #opensource. Retweets confirm that the recruitment messages reached open source participants on at least 5 continents (North America, South America, Europe, Africa, Australia) and a number of OSS developer communities (e.g. Python, Django, pan-African OSS, open science).
3.3. Validity & Reproducibility
@@
4. Results
We performed four analyses with the survey data. We began by examining how many participants work on different artifacts and use cases to confirm diverse labor participation in our study. Then, we examined how often participants receive compensation (specifically credit) and how visible their work is. Last, we considered the relationships between variables and calculated causal relationships where possible.
@@ 4.2. Half of OSS Labor May Not Receive Attribution
Overall, nearly 9 in 10 (89.1%) participants reported that some of their work is not fully visible (and therefore invisible) while only 1 in 10 (10.9%) reported full visibility (Fig. 1b).
@@ 4.3. Anchoring Shifts Perceptions of Labor Visibility
The group anchored to high visibility (shown in purple) viewed questions in descending order. They usually reported that $\geq 2$ people knew about their work “Sometimes” or “Often” over the past two years (Fig. 4b); then that 1 other person would “Often” or “Always” know about their work (Fig. 4c); and finally that “Rarely” or “Sometimes” would nobody else know about their work (Fig. 4d). Again, survey respondents tend to answer later survey questions by making adjustments from earlier questions (Tourangeau, 2018). This progression suggests the group’s relative perceptions of the labor visibilities; to them, their work was most likely to be partially visible (1 person), less likely to be fully visible ($\geq 2$ people), and less likely still to be not visible (0 people). Cumulatively, the cases that guarantee invisibility (when labor is either not visible or partially visible) is more likely than full visibility.

### v2

@@ 3. Methods
In this work, we applied a participatory feminist approach (Bardzell, 2010; Bardzell and Bardzell, 2011) to an explanatory mixed-methods design (Creswell and Clark, 2017). Our feminist approach identified an articulated need with members of several communities (practicing developer advocates [^6] The responsibilities of a developer advocate vary, but typically involve interfacing between an organization (which pays the individual) and users of the organization’s product. They do this by identifying user needs, advocating to the organization on behalf of those users, and advertising the organization’s products to current and potential users. and OSS community organizers); iteratively co-designed an approach for raising awareness of marginalized experiences in OSS development (both this study and a parallel series of workshops); and shared and discussed our results with members of these communities. Grounded in this approach, the explanatory sequential mixed-methods design (Creswell and Clark, 2017) followed a traditional approach of quantitatively answering a research question (RQ1), posing a new research question to explain the former results (RQ2), and answering the new question with qualitative analysis. The following sections describe our participatory process, how we designed and conducted the survey, and how we analyzed our data.
3.1. Participatory Process
To frame our participatory process, we begin with a positionality statement. As coauthors, we identify as English-speaking white academics and practitioners who reside in the West and Global North (United States and Australia). One of us identifies as non-binary, queer, and disabled; one as non-binary; one as a queer woman; and one as a cis man. Our experiences draw from computer science, engineering, and social science disciplines and epistemologies. These intersectional identities and experiences both broaden and constrain how we perceive and construct what it means to “do work” and study how people “do work” in OSS contexts.
Over about 26 months, we (all coauthors) held a series of biweekly meetings with one another, and intermittently with community members, on the topic of “OSS Ecosystems.” During the first 6 months, the topic of these meetings shifted toward “Ecosystem Mapping” motivated by a need among our developer advocates to characterize the diversity, quantity, interrelatedness, and vulnerability of tasks that took place throughout several global OSS communities. These meetings led to a series of 8 exploratory digital participatory mapping interviews inspired by physical participatory mapping activities (Chambers, 2006; Sletto, 2009; Zhou et al., 2016). The interviews were designed to understand the breadth, depth, and locales of activities that took place in support of four global Python language communities. In parallel, we conducted two pilot surveys about existing attribution tools that spanned multiple language communities and continents, their efficacies for documenting labor across their respective ecosystems, and developer satisfaction with different attribution techniques. The interviews and surveys led to a series of 6 multi-community workshops, organized by our team, with 33 community members from multiple communities across 5 continents (Africa, Asia, Australia, Europe, & North America). The workshops supported community members’ efforts “to identify the work which communities directly attribute to a specific outcome’s success” by helping community members define specific outcomes they invest in and identify “contributions [broadly defined] in a way that is explicit, granular, specific, and transparent.”
Our preliminary interviews, surveys, and workshops, while not the focus of this article, revealed to us that manual and automated tools have mixed records at measuring diverse activities, including both code and noncode activities, as others before us have shown (Monroy-Hernández et al., 2011; Young et al., 2021). This leaves many activities undocumented with repercussions for both individuals and organizations, as described above. Multiple community members expressed a need for better labor “taxonomies” and “metrics” to demonstrate how much work they do and what that work entails. To us, this resonated with D’Ignazio & Klein’s (D’Ignazio and Klein, 2020) argument that feminist data work should “make labor visible” to benefit marginalized computing workers. Our collective motivation for both the workshops and the design of this survey arose from the desire of community members to make their labor more visible when existing tools did not meet their needs. To that end, we saw need for a critical realist analysis (Terry et al., 2017) (acknowledging a plurality of sociocultural meanings and interpretations of labor and its visibility even as we sought to unearth an underlying “reality”) to measure the extent of OSS labor, its visibility, and invisibility. We hoped this would yield evidence that individuals of varied backgrounds could provide to traditionally positivist organizations about the extent of their labor.
We therefore co-designed a Qualtrics web survey (Tourangeau et al., 2013) to assess the pervasiveness of invisible labor. The two academic coauthors on our team drafted questions to measure labor focuses, plus the visibility and compensation dimensions of invisible labor. The full team of coauthors (academics, developer advocates, and community organizers) then iteratively revised all questions and their response options (particularly the artifact and use case categories), both among ourselves and with our Institutional Review Board (IRB), which approved the design before the start of the study. Because we were working with global OSS communities, and given our intent to release the raw data, our IRB recommended making our survey fully anonymous to ensure ethics compliance across multiple jurisdictions. This required us to avoid gathering any demographic data that could identify individuals given the small number of participants in certain demographic cross sections (e.g. an individual of a specific gender, project size, and specific country). The following sections detail the survey instrument’s design, distribution, and analysis.
After completing the initial survey analysis, we iteratively prepared a preprint of this article. We shared the preprint with community members to gauge how they felt about our findings, and they in turn shared it with other members of their communities. We plan to circulate this article as a more approachable presentation at developer conferences after peer review.
@@ 3.2. Survey Instrument Design
Our critical realist objectives (giving positivist organizations convincing representations of historically marginalized labor) informed our survey design process. To that end, the survey initially focused participant attention on recent collaborative OSS projects (still a more familiar and flexible term for participants than the research concept of an open source ecosystem), before broadening participant perception to the breadth of tasks performed in the ecosystem in which those projects were embedded. This gave us access to participant perceptions of tasks beyond code contributions while still including coding activities. We asked a series of questions about the “credit” they receive, which we defined for them as “the project told someone that you did something for the project,” as one form of compensation. Then, we asked several questions about how often other people know about the tasks they perform to measure labor visibility. See Appendix A for the full survey.
@@ 3.2.1. Guiding Participant Attention
The term “invisible labor” and topics related to supportive labor tend to be femininely and liberally coded (c.f. Frizzell, 2023; Kalita, 2023) — and even denigrated — in many global contexts (Berdahl et al., 2018; D’Ignazio and Klein, 2020). Socially coded contexts make it difficult to accurately capture lived experiences because participants tend to (a) self-select into or out of surveys and (b) signal through their responses based on how they want to be socially perceived, even if it does not represent their lived experience (Dell et al., 2012; Krumpal, 2013). Using coded language like “invisible labor” would have increased the likelihood of drawing ideologically aligned participants to the survey; dissuaded ideologically misaligned individuals from taking the survey; and elicited ideologically aligned responses from both groups. Therefore, our goal of accurately representing OSS community members’ lived experiences — regardless of their ideological leaning — led us to guide participant attention such that we limited gendered and political social signaling. Indeed, not limiting social desirability bias would have risked inaccurately representing all participants’ experiences in aggregate analyses in contradiction with our the critical realist objective of our feminist methodology: “making labor visible,” but in a way that positivist organizations also trust.
@@
A participant’s experience with a survey begins with recruitment materials. We therefore carefully avoided the term “invisible labor” in these materials, along with the consent form (Q1) and other questions throughout the survey. Instead, we identified another term common to open source contexts to help participants recall the full breadth of tasks that people perform, both inside formally-defined projects and outside them (that is, within the ecosystem). We advertised the survey around “receiving credit” for their work as this is an important concept to many in open source communities (c.f. Brand et al., 2015; Casari et al., 2021; Young et al., 2021) but which should be less likely to evoke either self-selection into the survey based on the survey focus or gendered social desirability bias (Krumpal, 2013). We acknowledge that no terminology is purely neutral nor uncoded, but believed “credit” was a common enough term in OSS (even if potentially masculinely coded) that it was less likely to invoke social signaling, insofar as possible while still asking about labor attribution.
@@
The survey began by prompting participants to focus on their experiences with open source projects that they worked on with other people during the last two years (Q2). Initially focusing their attention on collaborative open source projects gave participants the familiar idea of a “project” to anchor to (many would not have been familiar with “ecosystems” in this context) while giving us access to participant perceptions of OSS ecosystems that pose the greatest challenge to sustainability (National Science Foundation, 2023, 2024). Additionally, focusing their attention on collaborative contexts increased the likelihood that participants would more comprehensively consider their actions in the broader ecosystem (the collective organizing context) in which their project (the repository and traditionally related actions) is embedded. We limited the survey to a two-year time period to reduce experience underretrieval due to excessive cognitive load (Lenzner et al., 2010; Tourangeau, 2018).
@@
Later, we focused the participant’s attention on how “tasks vary in shape, size, and how visible they are to other people” (Q5). We asked them to “think through the different kinds of tasks you have done as part of these projects over the past 2 years.” We did this to provide rich retrieval cues about varied tasks and locales, which tend to improve recall (Tourangeau, 2018). This broadened their perception out from “just code” to include as many tasks as possible within their work, thereby bringing their visible and invisible labor within the broader ecosystem to the fore. For those who already considered non-coding activities to be contributions, this would not affect their judgment; however, for those who might consider code the only legitimate type of “contribution” (Casari et al., 2021), this design avoided the term contribution (via “task” instead) while improving information retrieval about diverse tasks (Tourangeau, 2018).
@@ 3.2.2. Compensation Questions
This followed with several questions about receiving credit for tasks (Q6-10). We asked about how many projects they received credit from (via 5-point Likert scale), how many tasks they received credit for (Likert scale), the medium through which they received credit (multiple select), and their satisfaction with the credit they received (Likert scale). These questions helped us gauge how much compensation (here, attribution) individuals receive, how they receive it, and to measure any relationship strength between attribution and perceptions of visibility.
@@ 3.2.3. Labor Visibility Questions
Unlike in crowd work where all work might take place on one platform (Toxtli et al., 2021), OSS activities take place across an ecosystem (Casari et al., 2021) that makes measurement via tracking difficult. We therefore measured labor visibility by asking participants how often different numbers of people knew about tasks they performed. This approximated how visible different tasks are to OSS communities by asking participants to implicitly assess the extent to which labor data is being created, shared, and made accessible. Combining randomization with human cognitive biases allowed us to estimate likely bounds on how much labor individuals perform at different visibility levels. [^7] Another alternative could have been to ask participants to identify all the tasks performed (or a representative sample of tasks); about the creation, sharing, and accessibility of data for each task; and to assess sharing and accessibility with respect to each person they interact. However, would induced significant cognitive load as each of these activities requires recalling information on the part of the participant, likely resulting in high drop-out rates. Instead, we opted for the method presented in the text. We began the labor visibility questions with a prompt (“Think about all the tasks you performed again,” Q11) to aid recall of their activities. Then, we presented participants with three questions:
@@
There are a few potential shortcomings to mitigate with this approach. For one, it may be more important for certain individuals to see some information than for others. However, this is often subjective, highly contextual, and difficult to estimate through a self-reported survey. Instead, we remain agnostic to the importance of information to specific individuals by asking about visibility to any participants at all with our method. For another, behavioral frequency impressions (e.g. how often someone performs invisible tasks) are prone to under- and overestimation (Conrad et al., 1998). While we cannot directly verify their perceptions, our approach treats participant perceptions as expertise (Bardzell, 2010). Third, information presented earlier in a survey tends to produce assimilation effects — that is, later judgments are moved in the direction of earlier judgments (Tourangeau, 2018), reminiscent of the anchoring and adjustment heuristic or priming effects (Kahneman et al., 1982; Molden, 2014). So if we only presented Q12, Q13, and Q14 in that order (called $\geq 2$, 1, 0 hereafter for clarity), participant responses would have been likely to bias results toward presumptions of high visibility (2 or more people, or “openness”). Therefore, we randomized whether participants saw these questions in order of what we call anchoring to high visibility ($\geq 2$, 1, 0) or anchoring to low visibility (0, 1, $\geq 2$).
@@
Following these, we asked participants how important it is to receive credit for tasks (Q16) along with several free response questions about what did and did not work well for them about how people give credit (Q17-19).
3.3. Survey Procedures
@@
To test the survey instrument and measure invisible labor in OSS, we recruited 142 participants between January and June, 2022. We did this by posting IRB-approved recruitment materials on the social media platform Twitter with the hashtag #opensource. Retweets confirm that the recruitment messages reached open source participants on at least 5 continents (North America, South America, Europe, Africa, & Australia) and a number of OSS developer communities (e.g. Python, Django, pan-African OSS, open science). Per our explanatory design, we answered RQ1 through four quantitative analyses and RQ2 through a subsequent thematic qualitative analysis.
Table 1. Example subset of participant responses to survey free response questions by participant ID with corresponding codes.
{tblr}
width = colspec = Q[70]Q[700]Q[130], hline1 = 0.5pt, row1 = Victoria,fg=white, row2 = Mischka, row6 = Mischka, row10 = Mischka, cell21 = c=30.943, cell61 = c=30.943, cell101 = c=30.943, Par. ID & Response Codes
Q17 - Think about how the projects you worked on gave people credit. From your perspective, what worked well about how they gave people credit?
P008 Release shout-outs and thanks are nice. announced,
genuine thanks
P025 The personal touch. An actual human commenting/noting a contribution is much more personally significant than an automated tool that merely collates automated metrics. manual, attributed, personal
P084 Git commit history for linux kernel is “permanent” people can always see your contribution *if they look for it*. attributed,
internal,
documented
Q18 - What did not work well about how they gave people credit?
P008 They misspelled my name on a thank-you plaque, that wasn’t fun. misattributed
P025 When credit mechanisms are arbitrary or completely automated. Completely automated tools don’t feel like your efforts are actually being noticed; if the recognition is arbitrary, it’s difficult to differentiate between ”this maintainer has a vendetta against me” and ”this maintainer forgot to recognise me”. opaque,
automated, unattributed
P084 So many people commit code to the linux kernel that no one bothers to look for your commits unless they need to. Until you are a “top 10” (or so) commiter to linux kernel, most people won’t know who you are or care.
And getting ”credit” for commits ”wears off” over time - so unless the commits were in the past year (for larger changes) or couple of months for smaller fixes, it doesn’t really matter.
internal, large, wears off
Q19 - Is there anything else that you’d like to share with us?
P008 (no response)
P025 My perception is that that projects (at least, the ones that have healthy communities) enthusiastically recognise contributions, and project maintainers are acutely aware of the way that a small recognition can matter to contributors.
However, larger organizations neither recognise nor appreciate the value of those contributions; and often have a very entitled attitude to what they can expect from (often volunteer) maintainers.
genuine thanks,
personal, large
P084 Credit for committing code is over rated. It’s so much easier to write new code than it is to debug someone else’s code (and post a fix) or review someone else’s code. People should be getting a lot more credit for reviewing code than committing code. It’s part of the reason being a subsystem or driver maintainer in the linux kernel matters way more than committing code - they don’t always get credit for this role though. certain tasks,
inconsistent,
attributed
3.3.1. Quantitative Analyses
Our first four analyses utilize descriptive statistics, correlation analysis, and linear regressions. We began by examining how many participants work on different artifacts and use cases to confirm diverse labor participation in our study. Next, we examined how often participants receive compensation (specifically credit) and how visible their work is. We then considered the relationships between variables and calculated causal relationships where possible.
@@ 3.3.1. Quantitative Analyses
3.3.2. Qualitative Analysis
Last, we conducted thematic analysis of qualitative responses to answer our second research question: What factors affect how commonly labor invisibility occurs in OSS ecosystems? We described textual responses through qualitative coding, iteratively defining themes that emerged from responses to questions about effective and ineffective credit practices. We followed a standard six step process for “Big Q” latent inductive thematic analysis (Braun and Clarke, 2012; Terry et al., 2017). A “Big Q” approach “operates within a qualitative paradigm and is characterized by (genuine) theoretical independence and flexibility, and organic processes of coding and theme development” (Terry et al., 2017, p.7). In other words, one researcher subjectively (but iteratively) familiarized themselves with the data, generated codes, constructed themes, revised themes, defined themes, and reported results (Terry et al., 2017). By “latent” coding we mean that our codes sought to capture implicit meanings within the text instead of merely capturing explicit meanings through “semantic” coding (Terry et al., 2017).
We began by reading the 162 qualitative responses to our free response questions: Q17 (69 responses to “what worked well?”), Q18 (63 responses to “what did not work well?”), and Q19 (30 responses to “anything else you’d like to share?”). Inductive coding initially yielded qualities of importance, and frustration, to participants. Some codes were rooted in participant language (e.g. that credit be “public,” ”individual,” “genuinely expressed”; and that it not be “inconsistent,” “arbitrary,” or “hard to find”) while other codes added slight abstraction (e.g. accessible, distributed, opaque, favoritism). Participant descriptions frequently contrasted with one another (some emphasize the necessity of “manual” attribution while others sought greater “automation,” for example). A second round of coding refined these initial codes, moving toward multidimensional descriptions of recurring concepts such as attribution (e.g. attributed, unattributed, misattributed) and motivation (e.g. career advancement, own-use, altruism, reputation). Table 1 shows examples of these 38 refined codes alongside corresponding participant responses. Recurring concepts gradually became our initial set of 6 themes (specifically: attribution/“giving credit,” in/accessibility, compensation, motivation, project characteristics, and process). Borrowing the techniques of asking questions and making comparisons from grounded theory (Corbin and Strauss, 2008), though, analytic memo writing led to a higher level of themes that both distinguish and relate these initial themes. We define and report these final themes toward the end of the Results section in answer to RQ2. But first, we turn to the results of our quantitative analyses in answer to RQ1.
4. Quantitative Evidence of Invisible Labor
Our quantitative analyses show that participants receive credit for roughly half of their work across diverse artifacts and use cases. Participants reported feeling only slightly satisfied with how often they receive credit and the mediums through which they receive it. We then found that visibility anchoring shapes perceptions of the visibility of a person’s work. The following sections detail these results.
@@ 4.2. Half of OSS Labor May Not Receive Attribution
Overall, nearly 9 in 10 (89.1%) participants reported that some of their work is not fully visible (and therefore invisible) while only 1 in 10 (10.9%) reported full visibility (Fig. 1b). While our finding on credit reception frequency cannot speak to other forms of compensation (e.g. pay, social status, opportunities), credit is one of the least financially and socially expensive forms of compensation suggesting that this compensation rate is potentially an upper bound on compensation frequency. Participants’ slight satisfaction with credit frequency and mediums may also suggest a possible explanation for why many individuals do not seek more credit. Many who were less satisfied with how much credit they receive listed sources of credit similar to those who were satisfied. However, satisfied individuals cited several sources that dissatisfied individuals cited less often including blogs (47.9% for satisfied, 5.3% for dissatisfied), presentations (37.5% vs. 10.5%), and project documentation (64.6% vs. 42.1%). It is possible, then, that these mediums might increase the efficacy of sharing information or the accessibility of information about OSS labor.
@@ 4.3. Anchoring Shifts Perceptions of Labor Visibility
The group anchored to high visibility (shown in purple) viewed questions in descending order. They usually reported that $\geq 2$ people knew about their work “Sometimes” or “Often” over the past two years (Fig. 4b); then that 1 other person would “Often” or “Always” know about their work (Fig. 4c); and finally that “Rarely” or “Sometimes” would nobody else know about their work (Fig. 4d). Again, survey respondents tend to answer later survey questions by making adjustments from earlier questions (Tourangeau, 2018). This progression suggests the group’s relative perceptions of the labor visibilities; to them, their work was most likely to be partially visible (1 person), less likely to be fully visible ($\geq 2$ people), and less likely still to be not visible (0 people). Cumulatively, the cases that satisfy the conditions for invisibility (when labor is either not visible or partially visible) are more likely than full visibility.